# WAFFLED: Exploiting Parsing Discrepancies to Bypass Web Application Firewalls

### v2

@@ 5.3. Web Application Frameworks Setup
5.4. Validation of Bypass Effectiveness
To validate WAF bypasses with malicious payloads, we employed a rigorous methodology. Each test began with the dispatch of an initial request, embedded with a malicious payload, through the WAF. Upon receipt by the web application framework, the framework attempted to parse the request and extract the content of the body, subsequently storing it in a designated file on the server.
Our validation process compared the original malicious payload with the content received and stored on the target server. A successful bypass was determined if these two elements – the initial payload and the stored content – were found to be identical. We developed a script that compared all input payloads with their corresponding stored versions on the server. This approach allowed us to precisely determine whether the malicious payload had circumvented the WAF and was accurately parsed by the target web framework.
This process was necessary because defining a successful bypass solely based on the ability of a request to evade the WAF would lead to misclassification. Such a definition could include malformed requests-those not conforming to HTTP standards or unprocessable by any framework-as successful bypasses, inflating false positives and obscuring real-world risks due to:
@@ 5.4. Validation of Bypass Effectiveness
Practical Mitigation Challenges. Malformed requests are frequently rejected at the framework level. Reporting these would inflating false positives and obscuring actionable findings.
Real-World Impact. A true attack occurs when a malicious request bypasses the WAF, is parsed by the framework, and the payload is executed. Malformed requests that frameworks cannot parse pose lower security risks and are lower priority.
All bypasses in our work represent successful attacks, meaning the request bypassed the WAF and was parsed successfully by a web application framework, enabling the malicious payload to be executed (e.g., XSS or SQL injection).
5.5. Experiment Results and Analysis
Table 2 provides the list of frameworks and Table 6 in the Appendix A provides the list of parsers that were inspected in our work. The experiments revealed that all of the frameworks and parsers were vulnerable to at least one bypass technique, highlighting the effectiveness and practical implications of our findings.
Table 2. List of tested web frameworks and whether they are vulnerable to at least one bypass in our study or not.
@@ 5.5. Experiment Results and Analysis
6. Findings
@@
6.1. Identification of Unique Bypass Requests
@@ 6.2. Classification of Bypasses
6.3. Bypassed Requests Analysis
@@
6.3.1. Multipart Bypass Classes
Our examination identified a total of 351 unique bypasses related to multipart content-type parsing. Figure 3 demonstrates valid bypass classes that we found in our work for multipart content-type. Table 3 contains examples for each discussed bypass category.
Figure 3. Occurrence of Successful Bypasses For Multipart Content-Types in Tested Frameworks
[…]
Table 3. Classification of Multipart Bypass Categories with Examples. Removals are displayed with a strike-through text with a gray background, while additions and replacements are shown with only a gray background.
@@ 6.3.1. Multipart Bypass Classes
6.3.2. XML Bypass Classes
Our analysis revealed a total of 299 unique bypasses for XML content types. Figure 4 demonstrates valid bypass classes for application/xml content-type. Table 4 contains examples for each discussed bypass category.
Figure 4. Occurrence of Successful Bypasses For application/xml Content-Types in Tested Frameworks
Extra Field Addition. This category involves adding an extra field outside the defined XML schema. For example, inserting a new field such as <field2>value2</field2> exploits schema validation weaknesses, allowing malicious payloads to bypass security checks.
DOCTYPE Closure Confusion. This class confuses the WAF by placing an extra character at the end of the xml body which confuses the WAF in parsing the DOCTYPE entity of the XML.
Schema Closure Manipulation. This technique manipulates the closure of XML schemas. By inserting characters, new elements, or duplicated field names and values at specific positions within the schema, the structure of the XML document is altered in a way that evades detection by the WAF.
@@ 6.3.2. XML Bypass Classes
Content-Type Header Replacement. In this technique, the parameter name within the Content-Type header is replaced with its value. This manipulation can confuse WAFs that rely on precise header structures for detection.
Misplaced Field. This method involves placing field values outside their corresponding XML tags. By altering the position of values within the XML document, the attack payload can bypass WAF rules that expect a specific structure.
Table 4. Classification of XML Bypass Categories with Examples
@@
6.3.3. JSON Bypass Classes
In our work, we identified a total of 557 unique bypasses for JSON content types. Figure 5 demonstrates valid bypass classes for application/json content-type. Table 5 contains examples for each discussed bypass category.
Figure 5. Occurrence of Successful Bypasses For application/json Content-Types in Tested Frameworks
@@ 6.3.3. JSON Bypass Classes
Table 5. Classification of JSON Bypass Categories with Examples.
@@ 6.4. Practicality of Findings
On each “forgot-password” page, we submitted the form after typing in a non-existent email address and intercepted the generated HTTP request on the Burp Suite tool. We examined the Content-Type header field and the request body format to decide the content-type for each request. We found that more than two-thirds of the websites use the application/x-www-form-urlencoded, while one-fourth of them use the application/json and its variation. Only two websites use multipart/form-data and one uses URL parameters.
For each intercepted “forgot-password” request, to check the interchangeability between content types, we used the “Change body encoding” feature of Burp Suite (in the “Repeater” section) to convert the content type to multipart/form-data in application/x-www-form-urlencoded requests. We then examined the response status code and response body to check the sameness of responses, which indicates that the application does not differentiate those two content types. We found that more than ninety percent of websites accept them both.
@@
7. Protecting WAFs and Frameworks by Normalizing HTTP Request Bodies
@@
7.1. Overview of HTTP Normalizer
The HTTP multipart body normalizer serves two primary functions: normalizing multipart request bodies, and rejecting non-compliant requests. Upon receiving an HTTP request, the normalizer separates the request’s headers and body using the AIOHTTP Python library. It then parses and validates the request’s Content-Type header, and if it indicates that the message body uses the MIME multipart encoding, it performs one of the following actions:
@@ 7.1. Overview of HTTP Normalizer
7.2. Normalizing Request Body
Currently, the normalizer applies only to requests that use a multipart transfer encoding. Future enhancements will extend this normalization capability to other content-types, aiming to demonstrate the viability of this approach in enhancing WAF security by rejecting malicious HTTP requests and reducing false positives through normalization. This methodology can be uniformly applied to all content types to achieve comprehensive normalization.
7.3. Evaluation of Normalizer
We evaluated the normalizer first by verifying that it does not reject any of a set of valid multipart message bodies. We obtained these from the test suites of the Tornado web server, RStudio, and the multiparty multipart message parser.
@@ 7.3. Evaluation of Normalizer
Table 6 demonstrates these evaluation metrics. Out of the 8 bypassable requests that are not rejected by the normalizer, all are blocked by Cloudflare’s WAF, resulting in a 100% success rate.
Table 6. Evaluation Metrics for HTTP Normalizer.
@@ 7.4. Performance Cost
The HTTP-Normalizer was developed to demonstrate that all identified bypass techniques can be prevented if WAFs adhere to proper parsing standards. Our findings, including the lack of successful bypasses against AWS WAF, validate this hypothesis. The intent of the project was not to create a fully optimized, production-grade tool with minimal overhead and broad content-type support, but rather to prove that an effective solution is feasible and implementable for a specific content type. By strictly adhering to RFC grammar rules, the HTTP-Normalizer effectively eliminates all bypass methods. While it currently functions as an additional evaluation layer, its rules can be seamlessly integrated into existing WAF rulesets with minimal performance impact.
8. Discussion
In this section, we discuss key considerations and challenges identified in our work, as well as exploring potential for future research and development based on our findings.
8.1. Potential for WAF and Framework Fingerprinting
Previous research has shown that it is possible to use HTTP parsing discrepancies to fingerprint web servers by utilizing a combination of requests known to trigger specific actions on target web servers (Topcuoglu et al., 2024). The findings of our work also show that parsing discrepancies are present in WAFs and web application frameworks, which opens up the possibility of fingerprinting them by sending requests known to bypass and those that do not bypass WAFs. A concrete example from our findings would be sending a request to the target from the JSON bypass class under the Field Name Manipulation category, which we know will only bypass and succeed if the target runs on Google Cloud Armor with the Spring Boot framework. We can send the request and then analyze the response to identify whether the underlying WAF is Google Cloud Armor. Future work could expand on this and propose a fingerprinting methodology on WAFs by using the findings of our work.
8.2. Practical Considerations
We focused on prominent web frameworks and their default or most popular parser packages. While multipart forms are common, not all frameworks have dedicated parsers for them, often leading to custom implementations from developers. These custom parsers can increase the risk of developer errors, potentially allowing bypassed requests to be parsed successfully server-side. However, our work did not examine custom parsers, leaving their implications outside the scope of our findings.
8.3. Limitations
This work primarily focuses on HTTP/1.1 web requests, thus future research could investigate HTTP/2. Also, the scope of WAFs tested may not cover all available systems. Future work could expand the range of tested frameworks and target additional content-types beyond those examined in this work. Exploring these areas would provide a more thorough understanding of WAF vulnerabilities and the potential for bypasses.
8.4. Usability vs. Security Debate
@@ 8.3. Limitations
While our findings show that WAFs can prevent these bypasses by adhering to proper RFC standards, real-world implementations may face challenges due to compatibility constraints or customer-specific requirements. Therefore, we do not blame WAF vendors for these bypasses, as they often face trade-offs between security and operational needs. However, the our proposed defense mechanism, the HTTP-Normalizer, demonstrates that all identified bypasses in this work are theoretically defendable, ensuring enhanced security without significantly impacting usability. Also, note that not all WAFs are susceptible to every type of bypass, suggesting that a well-designed parser adhering to all relevant RFC rules can mitigate these vulnerabilities.
9. Ethics Considerations and Disclosure Information
@@
All attacks and bypasses in this paper were conducted in a controlled environment using our infrastructure, ensuring no external impact. For our analysis, no malicious payloads were sent to the tested websites, and each site received no more than 10 requests.
All WAF tests adhered to bug bounty protocols, and bypasses were disclosed to the affected vendors, who subsequently acknowledged the existence of these vulnerabilities.
@@
  • Google Cloud Armor classified our report as a Tier 1, Priority 1, Severity 1 vulnerability under the ”Insecure by Default” category and rewarded us with a bug bounty.
@@
Our results and source code, including fuzzer input grammars, are publicly available [^4] https://github.com/sa-akhavani/waffled/. Sensitive bypass requests will remain restricted until the vulnerabilities are resolved.
10. Conclusion
@@
This paper’s investigation into popular WAFs, including Cloudflare, Cloud Armor, AWS WAF, Azure WAF, and ModSecurity, highlights significant concerns regarding parsing discrepancies in their request analysis. These discrepancies pose a substantial risk, as WAF users may falsely believe they are protected against common attack payloads.
As attackers employ increasingly sophisticated methods, the development of dynamic, intelligent WAFs becomes important. Our research not only contributes to the understanding of current WAF limitations, but also proposes effective solutions to defend against the identified threats. This work emphasizes the critical need for improved parsing consistency in WAFs to ensure robust protection against these attacks. Our proposed HTTP-Normalizer offers a promising solution, balancing security and usability by enforcing strict compliance with RFC standards.
@@ Acknowledgment
Appendix A Tested Parsers
Table 7. List of tested parsers, their weekly downloads as of July 2024 collected from NPM and PyPI Stats  [^6] https://www.npmjs.com and https://pypistats.orgwebsite, and whether they are vulnerable to at least one bypass in our study or not.
Parser Version Vulnerable Downloads
Node.js
Busboy 1.6.0 Yes 8,639,607
Formidable 3.5.1 Yes 6,998,735
fast-xml-parser 4.3.3 Yes 14,254,527
Python
xmltodict 0.13.0 Yes 10,374,671
xmlminidom 3.10.12 Yes n/a
python-multipart 0.0.9 Yes 4,288,034
Appendix B Web Framework Parsing Methods
Table 8 summarizes the methods used for parsing request bodies in our work in each framework. If a content-type is not supported by the default parser of a framework, popular third-party parsers are used instead.
Table 8. Request Body Parsing Methods in Tested Frameworks
Framework Content-Type Parsing Method
Express (Node.js) application/json request.body
FastAPI (Python) application/json request.json()
Laravel (PHP) application/json \$request->collect()
Spring Boot (Java) application/json Defined a class and used @RequestBody
Laravel (PHP) multipart/form-data \$request->collect()
Express (Busboy) multipart/form-data form.parse(request)
FastAPI (Python) multipart/form-data field1: str = Form("none")
Flask (Python) multipart/form-data request.form
Spring Boot (Java) multipart/form-data @RequestParam(value = "field1") String field1
Laravel (PHP) application/xml \$request->getContent()
Spring Boot (Java) application/xml Defined a class and used @RequestBody
Flask (xmltodict) application/xml parse(request.data)
Gin (Go) application/xml Defined an xmlform with field1 and field2

### v3

@@ 5.3 Web Application Frameworks Setup
5.4 Validation of Bypass Effectiveness
To validate WAF bypasses, we dispatched malicious requests through the WAF to target frameworks, which attempted to parse and store the body content on the server. We then compared the original payload with the stored result to determine if the bypass succeeded. A script automated this comparison, confirming whether the payload had both circumvented the WAF and been correctly parsed by the framework. This process was necessary because defining a successful bypass solely based on the ability of a request to evade the WAF would lead to misclassification. Such a definition could include malformed requests that are not conforming to HTTP standards or are unprocessable by any framework, as successful bypasses. This would lead to inflating false positives and obscuring real-world risks because of the following reasons:
@@ 5.4 Validation of Bypass Effectiveness
Practical Mitigation Challenges. Malformed requests are frequently rejected at the framework level. Reporting these would inflate false positives and obscure actionable findings.
A true attack occurs when a malicious request bypasses the WAF, is parsed by the framework, and the payload is executed. Malformed requests that frameworks cannot parse pose lower security risks and are lower priority. All bypasses in our work meet this criteria, enabling execution of the malicious payload (e.g., XSS or SQL injection).
5.5 Experiment Results and Analysis
Table II provides the list of frameworks and Table 4 provides the list of parsers that were inspected in our work. The experiments revealed that all of the frameworks and parsers were vulnerable to at least one bypass technique, highlighting the effectiveness and practical implications of our findings.
TABLE II: List of tested web frameworks and whether they are vulnerable to at least one bypass in our study or not.
@@ 5.5 Experiment Results and Analysis
TABLE III: List of tested parsers, their weekly downloads as of July 2024 collected from NPM and PyPI Stats  [^4] https://www.npmjs.com and https://pypistats.orgwebsite, and whether they are vulnerable to at least one bypass in our study or not.
Parser Version Vulnerable Downloads
Node.js
Busboy 1.6.0 Yes 8,639,607
Formidable 3.5.1 Yes 6,998,735
fast-xml-parser 4.3.3 Yes 14,254,527
Python
xmltodict 0.13.0 Yes 10,374,671
xmlminidom 3.10.12 Yes n/a
python-multipart 0.0.9 Yes 4,288,034
6 Findings
@@
6.1 Identification of Unique Bypass Requests
@@ 6.2 Classification of Bypasses
6.3 Bypassed Requests Analysis
@@
TABLE IV: Classification of Multipart Bypass Categories with Examples. Removals are displayed with a strike-through text with a gray background, while additions and replacements are shown with only a gray background.
@@
6.3.1 Multipart Bypass Classes
Our examination identified a total of 351 unique bypasses related to multipart content-type parsing. Figure 3 demonstrates valid bypass classes that we found in our work for multipart content-type. Table IV contains examples for each discussed bypass category.
Figure 3: Occurrence of Successful Bypasses For Multipart Content-Types in Tested Frameworks
[…]
6.3.2 XML Bypass Classes
Our analysis revealed a total of 299 unique bypasses for XML content types. Figure 4 demonstrates valid bypass classes for application/xml content-type. Table V contains examples for each discussed bypass category.
Figure 4: Occurrence of Successful Bypasses For application/xml Content-Types in Tested Frameworks
Extra Field Addition. Adding an extra field outside the defined XML schema. For example, inserting a new field such as <field2>value2</field2> exploits schema validation weaknesses, allowing malicious payloads to bypass security checks.
DOCTYPE Closure Confusion. Placing an extra character at the end of the xml body confuses the WAF in parsing the DOCTYPE entity of the XML.
Schema Closure Manipulation. This technique manipulates the closure of XML schemas. By inserting characters, new elements, or duplicated field names and values at specific positions within the schema, the structure of the XML document is altered in a way that evades detection.
@@ 6.3.2 XML Bypass Classes
Content-Type Header Replacement. The parameter name within the Content-Type header is replaced with its value. This manipulation can confuse WAFs that rely on precise header structures for detection.
Figure 5: Occurrence of Successful Bypasses For application/json Content-Types in Tested Frameworks
Misplaced Field. Placing field values outside their corresponding XML tags. By altering the position of values within the XML document, the attack payload can bypass WAF rules that expect a specific structure.
TABLE V: Classification of XML Bypass Categories with Examples
@@
6.3.3 JSON Bypass Classes
In our work, we identified a total of 557 unique bypasses for JSON content types. Figure 5 demonstrates valid bypass classes for application/json content-type. Table VI contains examples for each discussed bypass category.
@@ 6.3.3 JSON Bypass Classes
TABLE VI: Classification of JSON Bypass Categories with Examples.
@@ 6.4 Practicality of Findings
On each “forgot-password” page, we submitted the form after typing in a non-existent email address and intercepted the generated HTTP request on the Burp Suite tool. We examined the Content-Type header field and the request body format to decide the content-type for each request. We found that more than two-thirds of the websites use the application/x-www-form-urlencoded, while one-fourth of them use the application/json and its variation. Only two websites use multipart/form-data and one uses URL parameters. For each intercepted “forgot-password” request, to check the interchangeability between content types, we used the “Change body encoding” feature of Burp Suite (in the “Repeater” section) to convert the content type to multipart/form-data in application/x-www-form-urlencoded requests. We then examined the response status code and response body to check the sameness of responses, which indicates that the application does not differentiate those two content types. We found that more than ninety percent of websites accept them both.
@@
7 Protecting WAFs and Frameworks by Normalizing HTTP Request Bodies
@@
Overview of HTTP Normalizer. The HTTP multipart body normalizer serves two primary functions: normalizing multipart request bodies, and rejecting non-compliant requests. Upon receiving an HTTP request, the normalizer separates the request’s headers and body using the AIOHTTP Python library. It then parses and validates the request’s Content-Type header, and if it indicates that the message body uses the MIME multipart encoding, it performs one of the following actions:
@@
Normalizing Request Body. Currently, the normalizer applies only to requests that use a multipart transfer encoding. Future enhancements will extend this normalization capability to other content-types, aiming to demonstrate the viability of this approach in enhancing WAF security by rejecting malicious HTTP requests and reducing false positives through normalization. This methodology can be uniformly applied to all content types to achieve comprehensive normalization.
Evaluation of Normalizer. We evaluated the normalizer first by verifying that it does not reject any of a set of valid multipart message bodies. We obtained these from the test suites of the Tornado web server, RStudio, and the multiparty multipart message parser.
@@
Table VII demonstrates these evaluation metrics. Out of the 8 bypassable requests that are not rejected by the normalizer, all are blocked by Cloudflare’s WAF, resulting in a 100% success rate.
TABLE VII: Evaluation Metrics for HTTP Normalizer.
@@
Performance Cost. The HTTP-Normalizer was developed to demonstrate that all identified bypass techniques can be prevented if WAFs adhere to proper parsing standards. Our findings, including the lack of successful bypasses against AWS WAF, validate this hypothesis. The intent of the project was not to create a fully optimized, production-grade tool with minimal overhead and broad content-type support, but rather to prove that an effective solution is feasible and implementable for a specific content type. By strictly adhering to RFC grammar rules, the HTTP-Normalizer effectively eliminates all bypass methods. While it currently functions as an additional evaluation layer, its rules can be seamlessly integrated into existing WAF rulesets with minimal performance impact.
8 Discussion
@@
Usability vs. Security Debate. While WAFs can prevent these bypasses by following RFC standards, real-world deployments may face compatibility or customer-specific constraints. Thus, we do not blame vendors, as they often balance security with operational needs. However, our proposed defense mechanism, the HTTP-Normalizer, shows that all bypasses in this work are theoretically preventable without significantly impacting usability. Also, not all WAFs are vulnerable to every bypass in this research, indicating that a well-designed RFC-compliant parser can effectively mitigate such issues.
Practical Considerations. We focused on prominent web frameworks and their default or most popular parser packages. While multipart forms are common, not all frameworks have dedicated parsers for them, often leading to custom implementations from developers. These custom parsers can increase the risk of developer errors, potentially allowing bypassed requests to be parsed successfully server-side. However, our work did not examine custom parsers, leaving their implications outside the scope of our findings.
Limitations and Future Work. This work primarily focuses on HTTP/1.1 web requests, and future research could investigate bypass behaviors in HTTP/2. Additionally, while we covered major frameworks and WAFs, expanding the range of frameworks and content-types tested could provide broader insights. Our findings also highlight that parsing discrepancies can reveal fingerprinting opportunities for both WAFs and frameworks. Prior research has shown that HTTP parsing behavior can be used to fingerprint web servers [24], and our results show similar potential for WAFs. For example, sending a JSON bypass request under the Field Name Manipulation category can help identify Google Cloud Armor with the Spring Boot framework, as only that combination would parse and accept the payload. Future work could develop a systematic fingerprinting methodology based on these discrepancies.
9 Ethics and Disclosure Information
@@
All attacks and bypasses in this paper were conducted in a controlled environment using our infrastructure, ensuring no external impact. No malicious payloads were sent to the tested websites, and each site received no more than 10 requests. All WAF tests adhered to bug bounty protocols, and bypasses were disclosed to the affected vendors, who subsequently acknowledged the existence of these vulnerabilities.
@@
  • Google Cloud Armor classified our report as a Tier 1, Priority 1, Severity 2 vulnerability under the ”Insecure by Default” category and rewarded us with a bug bounty.
@@
10 Conclusion
@@
This paper’s investigation into popular WAFs, including Cloudflare, Cloud Armor, AWS WAF, Azure WAF, and ModSecurity, highlights significant concerns regarding parsing discrepancies in their request analysis. These discrepancies pose a substantial risk, as WAF users may falsely believe they are protected against common attack payloads. As attackers employ increasingly sophisticated methods, the development of dynamic, intelligent WAFs becomes important. Our research not only contributes to the understanding of current WAF limitations, but also proposes effective solutions to defend against the identified threats. This work emphasizes the critical need for improved parsing consistency in WAFs to ensure robust protection against these attacks. Our proposed HTTP-Normalizer offers a promising solution, balancing security and usability by enforcing strict compliance with RFC standards.